write decimal as a fraction: 3.4 the 4 has a line on top for repeating
Answer by Earlsdon(6287)

Write as a fraction: 3.444...
First, write it as a mixed number by changing the decimal part to a fraction.
Here's how you do that:
Change 0.444... to a fraction.
Let n = 0.444... Multiply both sides by 10
10n = 4.444... Now subtract n from 10n
10-n = 4.444... - 0.444... = 4, so:
10n-n = 9n = 4 Divide both sides by 9
9n/9 = 4/9
n = 4/9 But n = 0.444..., so
0.444... = 4/9
The mixed number is: 3 4/9 Now if you want this as an improper fraction, then:
3 4/9 = 31/9 Check this answer on your calculator.